Equipment and Calibration History
Page: 1 2 3 4 5 6 7 8 9 [10] 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 33 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70Ser No. | Date | Loc From | Title | Loc To | Notes |
---|---|---|---|---|---|
LI-7700 TG1-0421 | 2023-07-11 | Lab none | Installed | Hill Slough Gas Analyzer CH4 | Installed at Hill Slough to replace sn TG1-0211 for calibration. |
LI-7700 TG1-0211 | 2023-07-11 | Hill Slough Gas Analyzer CH4 | Removed | Lab none | Removed from Hill Slough for calibration. Replaced by sn TG1-0421. |
LI-7500A 75H-2514 | 2023-07-11 | Lab none | Installed | West Pond Gas Analyzer CO2/H2O | Installed at West Pond to replace sn 75H-2180 for calibration. |
LI-7500A 75H-2180 | 2023-07-11 | West Pond Gas Analyzer CO2/H2O | Removed | Lab none | Removed from West Pond for calibration. Replaced by sn 75H-2514. |
LI-7500A 75H-2667 | 2023-07-11 | Lab none | Installed | East End Gas Analyzer CO2/H2O | Installed at East End to replace sn 75H-2176 for calibration. |
LI-7500A 75H-2176 | 2023-07-11 | East End Gas Analyzer CO2/H2O | Removed | Lab none | Removed from East End for calibration. Replaced by sn 75H-2667. |
LI-7500A 75H-2182 | 2023-07-11 | Lab none | Installed | Mayberry Gas Analyzer CO2/H2O | Installed at Mayberry to replace sn 0418 for calibration. |
LI-7500A 0418 | 2023-07-11 | Mayberry Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Mayberry for calibration. Replaced by sn 75H-2182. |
CR1000 4190 | 2021-01-17 | Cal Kit Data Logger | Removed | Lab none | Removed from WP with the Cal Kit |
LI-7500A 0042 | 2023-06-13 | Lab | Calibration | Lab | Last calibration: 2022-11-29 AGC=100.3 SDM= DAC1= DAC2= AIU sn AIU-1569, Presure A0=58.004, A1=15.479 Ambient P=100.7kPa, Tair=19.8C, Tdew=11.0C Zero Air at 1150psi, AMP CA04119 524.77 at 1050 Tdewspan=16.01 CO2 zero 0.9478 -> 0.9486 Offset: 3.0ppm CO2 span 1.0145 -> 1.0184 Offset: -2.67ppm H2O zero 0.8472 -> 0.8461 Offset: -6.8ppm H2O span 1.0267 -> 1.039 Offset: -0.26ppm Notes: Calibrated with the IRGA vertical |
LI-7500A 0073 | 2023-06-13 | Lab | Calibration | Lab | Last calibration: 2022-11-15 AGC=102.2 SDM= DAC1= DAC2= AIU sn AIU-1569, Presure A0=58.004, A1=15.479 Ambient P=100.6kPa, Tair=18.8C, Tdew=11.8C Zero Air at 1150psi, AMP CA04119 524.77 at 1050 Tdewspan=16.01 CO2 zero 0.9165 -> 0.9161 Offset: -1.6ppm CO2 span 1.0085 -> 1.0093 Offset: -0.57ppm H2O zero 0.8296 -> 0.8328 Offset: 18.5ppm H2O span 1.0258 -> 1.0281 Offset: -0.4ppm Notes: Calibrate with IRGA vertical - slightly larger than usual H2O zero offset. |
LI-7500A 1594 | 2023-06-12 | Lab | Calibration | Lab | Last calibration: 2023-03-14 AGC=101.1 SDM=0 DAC1=- DAC2=- AIU sn AIU-1569, Presure A0=58.004, A1=15.479 Ambient P=100.6kPa, Tair=21.0C, Tdew=12.3C Zero Air at 1150psi, AMP CA04119 524.77 at 1050 Tdewspan=16.01 CO2 zero 0.8778 -> 0.8777 Offset: -0.6ppm CO2 span 1.0027 -> 1.0098 Offset: -5.07ppm H2O zero 0.8067 -> 0.8110 Offset: 25.5ppm H2O span 0.9989 -> 1.0126 Offset: -0.4ppm Notes: First lab calibration since repair at Licor - the usual larger offsets. Coefficients were all entered manually and the IRGA was vertical for this calibration. |
LI-7550 AIU-1870 | 2023-07-10 | Vaira none | Notes | Vaira none | Pressure coefficients found in a calibration file.
A0=58.383 A1=15.574 |
LI-7700 TG1-0421 | 2023-07-10 | Factory none | Calibration | Lab none | Received from the factory back at the lab. 13.8 hours of labor went into repairing the extensive water damage and instrument calibration. 18 parts were replaced. |
LI-7500A 1594 | 2023-06-16 | Lab none | Installed | Vaira Gas Analyzer CO2/H2O | Installed at Vaira to replace sn 75H-2514 for calibration. |
LI-7500A 75H-2514 | 2023-06-16 | Vaira Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Vaira for calibration. Replaced by sn 1594. |
LI-7500A 0073 | 2023-06-16 | Lab none | Installed | Tonzi-Floor none | Installed at Tonzi-Floor to replace sn 75H-2182 for calibration. |
LI-7500A 75H-2182 | 2023-06-16 | Tonzi-Floor Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Tonzi-Floor for calibration. Replaced by sn 0073. |
LI-7500A 0042 | 2023-06-16 | Lab none | Installed | Tonzi-Tower Gas Analyzer CO2/H2O | Installed at Tonzi-Tower to replace sn 75H-2667 for calibration. |
LI-7500A 75H-2667 | 2023-06-16 | Tonzi-Tower Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Tonzi-Tower for calibration. Replaced by sn 0042. |
CR23X 2303 | 2023-07-05 | Loaned none | Received | Lab none | Returned to the lab from Sophie Ruehr at the conclusion of her measurements along with PAR sensor sn 030382 |
PAR-LITE 030382 | 2023-07-05 | Loaned none | Received | Lab none | Returned to the lab from Sophie Ruehr at the conclusion of her measurements along with CR23X sn 2303 |
LI-7700 TG1-0211 | 2023-06-15 | Hill Slough | Calibration | Hill Slough | Calibration Location: Hill Slough
Last calibration: 2022-11-30 RSSI=73%, Optical RH=23% Ambient P=100.9kPa, Tair=17.0C Zero Air at 1200psi, AMP CA04119 4.2238 at 1000 CH4 Zero Offset: -.003ppm CH4 Span Offset: n/a Notes: Windy! Did not set the span because the sensor value would "stabilize" at different number that would vary by >0.25ppm. The "stable" value would change whenever a gust would blow through. |
LI-7700 TG1-0357 | 2023-06-14 | Bouldin Corn | Calibration | Bouldin Corn | Calibration Location: Bouldin Corn Last calibration: 2022-11-30 RSSI=75%, Optical RH=23% Ambient P=101.1kPa, Tair=23.0C Zero Air at 1250psi, AMP CA04119 4.2238 at 1125 CH4 Zero Offset: -.012ppm CH4 Span Offset: .0238ppm Notes: Breezy. Flushed headspace to 12% optics RH after calibration and swapped in a new desiccant bottle. |
LI-7700 TG1-0423 | 2023-06-15 | East End | Calibration | East End | Calibration Location: East End Last calibration: 2023-01-18 RSSI=82%, Optical RH=16% Ambient P=100.9kPa, Tair=26.7C Zero Air at 1150psi, AMP CA04119 4.2238 at 1000 CH4 Zero Offset: -.06ppm CH4 Span Offset: .0138ppm Notes: Optics RH=16%, which is pretty good, so we did not flush the headspace or change its desiccant. |
LI-7700 TG1-0155 | 2023-06-15 | West Pond | Calibration | West Pond | Calibration Location: West Pond Last calibration: 2023-01-18 RSSI=89%, Optical RH=13% Ambient P=101.0kPa, Tair=26.2C Zero Air at 1100psi, AMP CA04119 4.2238 at 1050 CH4 Zero Offset: 0ppm CH4 Span Offset: -.0088ppm Notes: Zero was right on! Since the optics RH=13% which is pretty good, we did not flush its headspace or change its desiccant. The heater and cleaning settings were set to the default (maybe because of recent power outages). I updated both to what I remember to be our usual settings. |
JAZ-EL350 JAZA2903 | 2023-06-13 | Loaned none | Returned | Lab none | Returned from Sophie in the past few weeks. Today I changed the clock from 16:39 to 15:30 to match the laptop time in PST. |
LI-7550 AIU-2312 | 2023-06-12 | Lab none | Returned | Disposed none | Returned to Andre Santos from Ameriflux. Returned with a Head 0128, power cable, the 7550 temperature sensor, ethernet cable, ethernet RJ45 adapter, and mounting plates. |
LI-7500A 75H-2418 | 2023-06-13 | Lab none | Returned | Disposed none | Returned to Andre Santos from Ameriflux. Returned with box AIU-1039, box brackets, head mount, IRGA cable and temp sensor. |
LI-7550 AIU-1039 | 2023-06-12 | Lab none | Returned | Disposed none | Returned to Andre Santos from Ameriflux. Returned with head 75H-2418, box brackets, head mount, IRGA cable and temp sensor. |
LI-7500A 1594 | 2023-05-30 | Factory none | Received | Lab none | Received back from Licor factory. Lamp was replaced and the sensor was recalibrated. |
LI-7500A 75H-2667 | 2023-04-14 | Lab none | Installed | Tonzi-Tower Gas Analyzer CO2/H2O | Installed at Tonzi to replace sn 1594, which probably needs to go back to the factory. |
LI-7500A 1594 | 2023-04-14 | Tonzi-Tower Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Tonzi-Tower for repair. It had zero signal strength when we arrived. Replaced by sn 75H-2667. |
MNKID-B KB04456 | 2023-02-15 | Lab none | Disposed | Disposed none | Given to Adina Paytan at UCSC to replace the controller that was flooded at their Porter site. They will give us a new one when the next order ships. |
LI-7550 AIU-1866 | 2023-05-17 | Lab none | Installed | Gilbert Tract Interface LI-7550 | Installed at GT after factory repair to replace sn AIU-1039, the AMP 7550. |
LI-7550 AIU-1039 | 2023-05-17 | Gilbert Tract Interface LI-7550 | Removed | Lab none | This AMP 7550 was removed from GT tower and replaced with out 7550 sn AIU-1866 which is back from the factory. |
eosFD FD20180026 | 2023-05-09 | Bouldin Corn Forced Diffusion C | Moved | Bouldin Alfalfa Forced Diffusion A | Moved from BC to BA to replace sn FD20160001 which broke. We should have soil CO2 flux data from Whendee's chamber measurements at BC. |
HFP01 11326 | 2023-05-03 | Bouldin Corn Heat Flux Plate A | Installed | Bouldin Corn Heat Flux Plate A | Installed in soil under tower at 2cm depth. |
HFP01 10867 | 2023-05-03 | Bouldin Corn Heat Flux Plate B | Installed | Bouldin Corn Heat Flux Plate B | Installed in soil under tower at 2cm depth. |
HFP01 10861 | 2023-05-03 | Bouldin Corn Heat Flux Plate C | Installed | Bouldin Corn Heat Flux Plate C | Installed in soil under tower at 2cm depth. |
CS655 17133 | 2023-05-03 | Bouldin Corn Soil Moisture Probe C | Installed | Bouldin Corn Soil Moisture Probe C | Installed in soil under tower. Probe was installed horizontally with top of probe 5cm under soil surface. |
CS655 17129 | 2023-05-03 | Bouldin Corn Soil Moisture Probe A | Installed | Bouldin Corn Soil Moisture Probe A | Installed in soil under tower. Probe was installed horizontally with top of probe 5cm under soil surface. |
CS655 17128 | 2023-05-03 | Bouldin Corn Soil Moisture Probe B | Installed | Bouldin Corn Soil Moisture Probe B | Installed in soil under tower. Probe was installed horizontally with top of probe 5cm under soil surface. |
HFP01 11326 | 2023-04-26 | Bouldin Corn Heat Flux Plate A | Moved | Bouldin Corn Heat Flux Plate A | Moved from field to tower in preparation for impending plowing. Not installed in soil yet. |
HFP01 10867 | 2023-04-26 | Bouldin Corn Heat Flux Plate B | Moved | Bouldin Corn Heat Flux Plate B | Moved from field to tower in preparation for impending plowing. Not installed in soil yet. |
HFP01 10861 | 2023-04-26 | Bouldin Corn Heat Flux Plate C | Moved | Bouldin Corn Heat Flux Plate C | Moved from field to tower in preparation for impending plowing. Not installed in soil yet. |
CS655 17133 | 2023-04-26 | Bouldin Corn Soil Moisture Probe C | Moved | Bouldin Corn Soil Moisture Probe C | Moved from field to tower in preparation for impending plowing. Not installed in soil yet. |
CS655 17129 | 2023-04-26 | Bouldin Corn Soil Moisture Probe A | Moved | Bouldin Corn Soil Moisture Probe A | Moved from field to tower in preparation for impending plowing. Not installed in soil yet. |
CS655 17128 | 2023-04-26 | Bouldin Corn Soil Moisture Probe B | Moved | Bouldin Corn Soil Moisture Probe B | Moved from field to tower in preparation for impending plowing. Not installed in soil yet. |
EXO2 22C103416 | 2023-04-25 | Hill Slough | Removed | Lab none | Removed from HS channel for routine calibration |
EXO pH-ORP 22D102641 | 2023-04-25 | Hill Slough pH-ORP | Removed | Lab none | Removed from HS channel for routine calibration |
EXO fDOM 22E100649 | 2023-04-25 | Hill Slough fDOM | Removed | Lab none | Removed from HS channel for routine calibration |
EXO Turbidity 22E106072 | 2023-04-25 | Hill Slough Turbidity | Removed | Lab none | Removed from HS channel for routine calibration |
LI-7500A 0041 | 2023-04-25 | Lab none | Installed | Hill Slough Gas Analyzer CO2/H2O | Installed at Hill Slough to replace sn 0042 for calibration. |
LI-7500A 0042 | 2023-04-25 | Hill Slough Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Hill Slough for calibration. Replaced by sn 0041. |
HMP155 T2930114 | 2023-04-26 | Lab none | Installed | Cal Kit RH & Temp | Installed as part of the calkit at Tonzi Floor |
CR1000 4190 | 2023-04-26 | Lab none | Installed | Cal Kit Data Logger | Installed as part of the calkit at TZ Floor |
LI-7550 AIU-1866 | 2023-04-27 | Factory none | Received | Lab none | Received back from Licor after repair. The main power supply was found defective and replaced. |
LI-7500A 0041 | 2023-04-24 | Lab | Calibration | Lab | Last calibration: 2023-01-05 AGC=100 SDM= DAC1= DAC2= Ambient P=100.4kPa, Tair=21.7C, Tdew=6.85C Zero Air at 1300psi, AMP CA04119 524.77 at 1150 Tdewspan=16.00 CO2 zero 0.9961 -> .9968 Offset: 2.54ppm CO2 span 1.0234 -> 1.0202 Offset: 2.12ppm H2O zero 0.8232 -> .8222 Offset: -5.8ppm H2O span 1.0382 -> 1.0302 Offset: .17ppm Notes: Span point entered as "524.67" for sensor to read ~524.77 after spanning |
eosFD 20160002 | 2018-01-18 | Factory none | Disposed | Disposed none | Marked as disposed |
eosFD 20160003 | 2018-01-18 | Factory none | Disposed | Disposed none | Marked as disposed |
eosFD 20160004 | 2018-01-18 | Factory none | Disposed | Disposed none | Marked as disposed |
eosFD 20160005 | 2018-01-18 | Factory none | Disposed | Disposed none | Marked as disposed |
eosFD 20160006 | 2018-01-18 | Sherman Barn none | Disposed | Disposed none | Marked as disposed |
eosFD FD20180068 | 2022-08-24 | Factory none | Disposed | Disposed none | This unit was originally borrowed from Eosense. Since this model is no longer in production they can not repair it ahead of other purchased units so we will not be getting it back. |
eosFD FD20180068 | 2022-08-23 | Lab none | Shipped | Factory none | Sent to Eosense on RMA #220809 for repair. |
eosFD FD20190028 | 2022-08-24 | Factory none | Disposed | Disposed none | This unit was originally borrowed from Eosense. Since this model is no longer in production they can not repair it ahead of other purchased units so we will not be getting it back. |
eosFD FD20190028 | 2022-08-23 | Lab none | Shipped | Factory none | Sent to Eosense for evaluation and repair. |
LI-7500A 0035 | 2023-04-11 | Lab none | Installed | Bouldin Alfalfa Gas Analyzer CO2/H2O | Installed at Bouldin Alfalfa to replace sn 0073 for calibration. |
LI-7500A 0073 | 2023-04-11 | Bouldin Alfalfa Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Bouldin Alfalfa for calibration. Replaced by sn 0035. |
eosFD FD20160001 | 2023-04-11 | Bouldin Alfalfa Forced Diffusion A | Removed | Lab none | Removed from BA underneath the tower. The data has been super noisy since mid March. The cable is still on the tower. |
LI-7500A 75H-2669 | 2023-04-11 | Lab none | Installed | Bouldin Corn Gas Analyzer CO2/H2O | Installed at Bouldin Corn to replace sn 0041 for calibration. |
LI-7500A 0041 | 2023-04-11 | Bouldin Corn Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Bouldin Corn for calibration. Replaced by sn 75H-2669. |
HygroVUE10 E3514 | 2023-04-11 | Lab none | Installed | Bouldin Corn RH & Temp | Installed at BC tower at the same height/location to replace RH/T sensor HMP60 sn L1740398. |
LI-7500A 0065 | 2023-04-11 | Lab none | Installed | Gilbert Tract Gas Analyzer CO2/H2O | Installed at Gilbert Tract to replace sn 75H-2667 for calibration. |
LI-7500A 75H-2667 | 2023-04-11 | Gilbert Tract Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Gilbert Tract for calibration. Replaced by sn 0065. |
LI-7500A 75H-2667 | 2023-04-13 | Lab | Calibration | Lab | Last calibration: 2023-01-13 AGC=101 SDM=0 DAC1=- DAC2=- Ambient P=100.2kPa, Tair=23.1C, Tdew=-1.99C Zero Air at 1200psi, AMP CA04119 524.77 at 1150 Tdewspan=16.0 CO2 zero 0.9418 -> 0.9417 Offset: -0.2ppm CO2 span 1.0233 -> 1.0213 Offset: 1.43ppm H2O zero 0.8421 -> 0.8141 Offset: -4.1ppm H2O span 1.0314 -> 1.0317 Offset: 0.0ppm Notes: |
LI-7500A 0065 | 2023-04-10 | Lab | Calibration | Lab | Last calibration: 2023-01-04 AGC=102 SDM= DAC1= DAC2= Ambient P=100.6kPa, Tair=23.7C, Tdew=10.5C Zero Air at 1300psi, AMP CA04119 524.77 at 1150 Tdewspan=16.00 CO2 zero 0.8915 -> .8918 Offset: 1.72ppm CO2 span 1.0227 -> 1.0217 Offset: .83ppm H2O zero 0.7323 -> .7295 Offset: -20.4ppm H2O span 1.0226 -> 1.0128 Offset: .21ppm Notes: |
LI-7500A 0035 | 2023-04-10 | Lab none | Calibration | Lab none | Last calibration: 2023-01-13
AGC=99.3 SDM= DAC1= DAC2= Ambient P=100.6kPa, Tair=21.0C, Tdew=10.5C Zero Air at 1400psi, AMP CA04119 524.77 at 1150 Tdewspan=16.00 CO2 zero 0.9374 -> .9375 Offset: .28ppm CO2 span 1.0295 -> 1.0253 Offset: 2.83ppm H2O zero 0.9003 -> .8994 Offset: -5.1ppm H2O span 1.0369 -> 1.0291 Offset: .16ppm Notes: Used "Span" point of 524.67 to get sensor to read 524.77ppm after span |
LI-7500A 75H-2669 | 2023-04-07 | Lab none | Maintenance | Lab none | Chemical swap. The magnesium perchlorate was a white powdery mass and had to be loosened with a screwdriver before it would all come out. The Ascarite II came out easily. I had tried to calibrate the sensor eariler in the day but the H2O zero offset was -38.3 mmol/m3 and took a long time (45+min) to stabilize. The H2O span never stabilized. |
LI-7500A 75H-2669 | 2023-04-10 | Lab | Calibration | Lab | Last calibration: 2023-01-04 AGC=101.5 SDM= DAC1= DAC2= Ambient P=100.6kPa, Tair=24.9C, Tdew=10.8C Zero Air at 1300psi, AMP CA04119 524.77 at 1150 Tdewspan=16.00 CO2 zero 0.9187 -> .9207 Offset: 7.69ppm CO2 span 1.0187 -> 1.0176 Offset: 0.83ppm H2O zero 0.9029 -> .9001 Offset: -17.3ppm H2O span 1.033 -> 1.0105 Offset: .47ppm Notes: Offsets are recorded using the 2023-01-04 calibration as the starting config file. Attempted calibration on 4/7 but the H2O zero offset was -38.3 after needing a long time (45+min) to stabilize. The H2O span offset never stabilized. I decided to swap the chemicals then and this is the first calibration since chemical swap. |
JAZ-EL350 JAZA2903 | 2023-04-07 | Lab none | Loaned | Loaned none | Lent to Sophie Ruehr for measuring plants/paint chips (trying to find a non-NIR-reflective paint to make a high-NIR-contrast photo background for taking SIF pictures of plants) |
eosFD FD20180026 | 2023-03-08 | Lab none | Installed | Bouldin Corn Forced Diffusion C | Installed under BC tower to replace sn FD20180067 which got flooded |
LI-7550 AIU-0674 | 2023-03-08 | Lab none | Installed | Bouldin Corn Interface LI-7550 | Installed at BC tower to replace sn AIU-1866 which was removed because of power problems |
HMP60 L1740398 | 2023-04-03 | Bouldin Corn RH & Temp | Removed | Lab none | Removed from the BC tower and brought back to the lab for testing b/c the sensor was not giving any data |
LI-7500A 75H-2176 | 2023-04-03 | Lab none | Installed | East End Gas Analyzer CO2/H2O | Installed at East End to replace sn 75H-2669 for calibration. |
LI-7500A 75H-2669 | 2023-04-03 | East End Gas Analyzer CO2/H2O | Removed | Lab none | Removed from East End for calibration. Replaced by sn 75H-2176. |
LI-7500A 75H-2180 | 2023-04-03 | Lab none | Installed | West Pond Gas Analyzer CO2/H2O | Installed at West Pond to replace sn 0065 for calibration. |
LI-7500A 0065 | 2023-04-03 | West Pond Gas Analyzer CO2/H2O | Removed | Lab none | Removed from West Pond for calibration. Replaced by sn 75H-2180. |
LI-7500A 0418 | 2023-04-03 | Lab none | Installed | Mayberry Gas Analyzer CO2/H2O | Installed at Mayberry to replace sn 0035 for calibration. |
LI-7500A 0035 | 2023-04-03 | Mayberry Gas Analyzer CO2/H2O | Removed | Lab none | Removed from Mayberry for calibration. Replaced by sn 0418. |
JAZ-EL350 JAZA2903 | 2023-04-06 | Lab none | Adjusted | Lab none | Adjusted clock from 15:03 to 15:45 PST. This clock probably needs to be reset every few months. |
HMP45A S4220026 | 2023-03-29 | Lab none | Serviced | Lab none | Cut off the 75' cable to solder on to HMP45 sn D4420029, which will be used by Allen Goldstein's lab on their BWW urban tower. The sensor itself is still stashed in the sensor cabinet because it might have some useable parts. |
HMP45AC D4420029 | 2023-03-30 | Lab none | Loaned | Loaned none | Given to Erin Katz from Allen Goldstein's lab. They will use this sensor on their BWW urban tower. |
HMP45AC D4420029 | 2023-03-29 | Lab none | Serviced | Lab none | Spliced on 75' cable from HMP45A sn S4220026 so the new cable length is 125'. I ran this sensor for a few hours in the lab and all seemed well. This sensor will be used by Allen Goldstein's lab on their BWW urban tower. The original HMP on the urban tower was connected to the analog inputs on the 7700 sn TG1-0421, but we've had to remove the 7700 because of water damage so they need an HMP with a long cable to hook up to their boiler room equipment. |
LI-7700 TG1-0421 | 2023-03-28 | Lab none | Shipped | Factory none | Shipped back to Licor for repair. Expecting a 4-6 months turnaround time |
LI-7700 TG1-0421 | 2023-03-24 | Loaned none | Removed | Lab none | Removed from the BWW urban tower because it was not working. Joe opened it and found a bunch of water inside the bottom can. There is corrosion on the boards and will need to be returned to Licor for repair. |
LI-7550 AIU-1866 | 2023-03-03 | Lab none | Shipped | Factory none | Shipped to Licor for repair |
CR23X 2303 | 2023-04-05 | Lab none | Loaned | Loaned none | Loaned to Sophie Ruehr with PAR sn 030382 for field measurements (light extinction curves at the Lake Merrit bonsai garden) |
PAR-LITE 030382 | 2023-04-05 | Lab none | Loaned | Loaned none | Loaned to Sophie Ruehr with 23x sn 2303 for field measurements (light extinction curves at the Lake Merrit bonsai garden) |